OpenStudy (anonymous):

Check my answer? x-6/x<=-1

OpenStudy (anonymous):

I got x<=-3 or x>=0

OpenStudy (anonymous):

\[x-\frac{ 6 }{ x }\le-1\]

OpenStudy (anonymous):

\[x-\frac{ 6 }{ x }+1\le0\]

OpenStudy (anonymous):

x>0? @AliLnn

OpenStudy (anonymous):

Go on cymath.com and you will find your answer